“Wonderful! Not only did our salespeople do a good job in meeting the sales budget this year, but our production people did a good job in controlling costs as well,” said Kim Clark, president of Martell Company. “Our $18,300 overall manufacturing cost variance is only 1.2% of the $1,525,000 standard cost of products made during the year. That’s well within the 3% parameter set by management for acceptable variances. It looks like everyone will be in line for a bonus this year.”

The company produces and sells a single product. The standard cost card for the product follows:

Inputs (1)
Standard
Quantity
or Hours
(2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2 feet $ 8.45 per foot $ 16.90
Direct labor 1.4 hours $ 16 per hour 22.40
Variable overhead 1.4 hours $ 2.50 per hour 3.50
Fixed overhead 1.4 hours $ 6 per hour 8.40
Total standard cost per unit $ 51.20

The following additional information is available for the year just completed:

  1. The company manufactured 30,000 units of product during the year.
  2. A total of 64,000 feet of material was purchased during the year at a cost of $8.55 per foot. All of this material was used to manufacture the 30,000 units produced. There were no beginning or ending inventories for the year.
  3. The company worked 43,500 direct labor-hours during the year at a direct labor cost of $15.80 per hour.
  4. Overhead is applied to products on the basis of standard direct labor-hours. Data relating to manufacturing overhead costs follow:
Denominator activity level (direct labor-hours) 35,000
Budgeted fixed overhead costs $ 210,000
Actual variable overhead costs incurred $ 108,000
Actual fixed overhead costs incurred $ 211,800

Required:

1. Compute the materials price and quantity variances for the year.

2. Compute the labor rate and efficiency variances for the year.

3. For manufacturing overhead compute:

a. The variable overhead rate and efficiency variances for the year.

b. The fixed overhead budget and volume variances for the year.
Answer #1
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($8.45 - $8.55) × 64000 = $6400 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((2Feet X 30000Unit)-64000 Unit )X $8.45 = $33800U F
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($16/hour - $15.80/hour) × 43500 Hours= $8700 F
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (1.4 Hour × 30000 Unit - 43500) × $16 = $24000 UF
Calculation of Variable OH rate variance
= (Standard Variable OH per hour - actual variable OH per hour) × Actual hours used
= ($2.5/hour - $2.48/hour) × 43500 Hours= $870 F
* Actual Cost /Hour= $108000/43500=$2.48/Hour
Calculation of Variale OH efficiency variance:
= (standard hours required for actual production - actual hours used) × standard Rate
= (1.4 Hour × 30000 Unit -43500 ) × $2.5= $3750 U F
Fixed OH Spending Variannce = Budgeted   OH- Actul OH
210000-$108000=$102000 F
Fixed OH Volume Variannce = Applied    OH- Budgeted OH
((1.4 Hour X30000 X6)-210000=$42000F